# Finite Difference Solution Ansatz approach in Least-Squares Monte Carlo

### v9

@@ Abstract
This article presents a simple but effective and efficient approach to improve the accuracy and stability of Least-Squares Monte Carlo. The key idea is to construct the ansatz of conditional expected continuation payoff using the finite difference solution from one dimension, to be used in linear regression. This approach bridges between solving backward partial differential equations and Monte Carlo simulation, aiming at achieving the best of both worlds. In a general setting encompassing both local and stochastic volatility models, the ansatz is proven to act as a control variate, reducing the mean squared error, thereby leading to a reduction of the final pricing error. We illustrate the technique with realistic examples including Bermudan options, worst of issuer callable notes and expected positive exposure on European options under valuation adjustments.
Keywords: Least-squares regression, finite difference method, partial differential equations, Bermudan option, worst of issuer callable note, credit valuation adjustment, high-dimensional derivative pricing
@@ 1 Introduction
Solving partial differential equations (PDE) with finite difference methods and Monte Carlo simulation are two major advanced numerical methods for exotic derivative pricing among practitioners. As the dimensionality of the problem increases, the PDE approach becomes less feasible due to implementation complexity and the numerical burden; specifically, there is no efficient and accurate implementation widely accepted in the industry with dimension greater than two. On the other hand, the Monte Carlo method is widely used in high-dimensional derivative pricing without suffering from the ‘curse of dimensionality’ as PDE with grid approach does.
@@
In this paper, we propose an enhanced least-squares regression based method by Longstaff and Schwartz, (2001), which aims at improving the projection of conditional expectations by constructing an ansatz from a finite difference PDE solver, to be used in the linear regression. Previous studies in the literature have explored methodologies that aim to combine PDE-based and simulation-based approaches, to leverage the strengths of both. For instance, Farahany et al., (2020) introduced a method by mixing a Fourier transforms based PDE solver on discretized volatility space and regressing the value function onto these resulting conditional payoff profiles, under stochastic volatility model. However, our work is distinguished from their proposal in below aspects:
@@
The rest of this paper is organized as follows. Sec.2 is the description of the framework from a theoretical standpoint, where the underlying model settings are presented in Sec.2.1. Then we introduce background knowledge in Sec.2.2 in the context of American-style exercise derivatives pricing and CVA. After setting up the general problem in Sec.2.2.1, we outline prevailing numerical methods in practice, including the least-squares Monte Carlo approach (Sec.2.2.2 and its limitation in Sec.2.2.3), and finite difference PDE method (Sec.2.2.4). With such foundation, and further motivated by the analysis of a limiting case in Sec.2.3.1, we proceed to introduce our novel approach in Sec.2.3, by highlighting its enhancement over the classical approach in Sec.2.3.2 and subsequently performing error analysis in Sec.2.3.3. Then implementation details will be given in Sec.2.3.4. The numerical results are delivered in Section 3, which compares the new method against existing alternatives regarding accuracy, efficiency, and stability: Sec.3.1, 3.2, and 3.3 examine Bermudan option pricing under the Black-Scholes (up to 4D) and Heston models; Sec.3.4 details CVA calculations for European options under 4D; last but not least, Sec.3.5 presents examples of worst-of issuer callable notes under a multi-variate local volatility model. Finally, concluding remarks are given in Sec.4.
@@ 2.2.1 Problem Formulation
As a general framework relevant for derivative valuation as well as regulatory CVA computations, we assume that throughout a given terminal expiry $T$ with the riskless interest rate $r$ defined previously, there exists a complete probability space ($\Omega,\operatorname{\mathcal{F}},(\operatorname{\mathcal{F}}_{t})_{0\leq t\leq T},\operatorname*{\mathbb{Q}}$), accommodating the model settings in Sec.2.1, under risk-neutral measure $\operatorname*{\mathbb{Q}}$. Also, let $L^{2}(\Omega,\operatorname{\mathcal{F}},(\operatorname{\mathcal{F}}_{t})_{0\leq t\leq T},\operatorname*{\mathbb{Q}})$, hereinafter simplified as $L^{2}$, denote the space of square integrable functions. As problem formulation, from $t=0$ and final expiry $T$ of an derivative contract, we define a set of discretized time points as $\pi(t):=\{t_{k}\geq t|t_{k}=k\Delta T,1\leq k\leq M=\frac{T}{\Delta T}\}$ with a time interval $\Delta T$. $\pi(t)$ represents either the exercise times for American-style option, or discrete exposure monitoring times in CVA calculation. Within the complete probability space $\Omega$, there exists a discrete filtration $\{\operatorname{\mathcal{F}}_{t_{k}}|1\leq k\leq M\}$ on $\pi(0)$. For notational clarity, we use $XYZ(\omega,t)$ to denote a stochastic variable under $\operatorname{\mathcal{F}}_{t}$ with $\omega\in\Omega$. In this spirit, the state variables are denoted by $X(\omega,t_{k})$, which is $\operatorname{\mathcal{F}}_{t_{k}}$-measurable.
@@
$PV(CVA)=\operatorname{\mathbb{E}}[\Phi^{PV(CVA)}(\{Z(\omega,t_{k}),\tilde{V}(\omega,\tilde{\tau}(t_{k})),\operatorname{\mathbb{E}}[Y^{PV(CVA)}(\omega,t)|\operatorname{\mathcal{F}}_{t_{k}}]|t_{k}\in\pi(0)\})|\operatorname{\mathcal{F}}_{0}].$ (2.3)
@@ 2.2.2 Least-Squares Monte Carlo
In Eqn.2.5, $\mathcal{L}(Y(\omega,t);\vec{b};\vec{\phi}^{P}_{R}(X(\omega,t_{k})))$ is the objective function to be minimized, which is the residual sum of squares (RSS). Its lower bound is given by, in the limit of $\vec{b}\cdot\vec{\phi}^{P}_{R}(X(\omega,t_{k}))\to\operatorname{\mathbb{E}}[Y(\omega,t)|X(\omega,t_{k})]$,
@@
where the exercise indicator function is defined as $I(\omega,t_{k}):=\Theta\left(\pm[Z(\omega,t_{k})-F(\omega,t_{k})]\right)$, with Heaviside step function denoted by $\Theta(x)$ and ‘$+$’(‘--’) for holder (issuer) exercise. The formal dynamic programming solution of $V(\omega,\tau(t_{k}))$ is also given by
@@
[…]
@@
$\begin{split}\widehat{CVA}=\frac{1-R_{ctpy}}{N_{p}}\sum_{i=1}^{N_{p}}\sum_{k=1}^{M}\Phi^{CVA(i)}_{t_{k}},\end{split}$ (2.19)
@@ 2.2.3 Basis function and Limitations
2.2.4 Finite difference approach in PDE
As opposed to LSM, alternative numerical approach to solve the American-style option pricing is the finite difference (FD) method. The elegance of this method lies in the fact that it marches from maturity to present, which naturally fits into the requirement of backward recursive optimization in stopping time problems. We restrict the discussion within 1D for maximum efficiency. The local volatility setting as describe in Sec.2.1 with $d=1$ is solvable under 1D FD PDE, under which the option price $V^{FD}(S,t)$ as function of underlying spot $S$ (omitting the subscript $i$ for the mono underlying case) and time $t$ satisfies:
@@ 2.3.2 FD Solution: Control Variate
Here, we decompose an $L^{2}$ projection using two sets of basis functions, $\vec{\phi}^{P^{\text{FD}}}(x)$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$, effectively implementing a Frisch-Waugh-Lovell (FWL) type decomposition into and $P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}$ and $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$. We will focus on a detailed examination of $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$, noting that $P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}$ can be analyzed symmetrically by swapping the roles of the FD and LSM basis functions.
@@ 2.3.3 Error Analysis
$\begin{split}M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}]&=M^{\text{FD}}_{k}[P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})]]\\ &=\vec{\beta}^{P^{\text{LSM}\perp{\text{FD}}}}_{k}(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D}))\cdot M^{\text{FD}}_{k}[\vec{\phi}^{P^{\text{LSM}}}_{R-1,k}]\\ &=(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}T}_{R-1,k}])^{-1}\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})]\cdot\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}.\\ \end{split}$ (2.40)
@@
  • In the limit of $\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})\to 1$, the result degenerates to Lemma 1 as expected.
  • In the limit of $\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})\to 0$, the control variate has null effect as $f^{\text{FD}}_{k}(X_{k})$ is orthogonal to $F_{k}$ in $L^{2}$. In this regime, the remaining mild error reduction, going as $\sqrt{\frac{R-1}{R}}$ is purely from reducing susceptibility of overfitting by removing unnecessary high order monomial component under the constraint of Eqn.2.38; further taking the limit of $R\to\infty$, $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ converges to $P_{k}^{\text{LSM}}$ with MSE being equal to each other.
@@
in the limit of $M\to\infty$, where $\mathcal{X}^{P}_{0}:=\lim_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}\mathcal{X}^{P}_{k}$ and $\mathcal{Y}^{P}_{0}:=\lim_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}\mathcal{Y}^{P}_{k}$. For avoidance of doubt, we use “0” as the subscript to represent this limiting case where $k$ is indexed from 1 as $k\in[1,M]$.
@@
$\begin{split}\operatorname{\mathbb{E}}(V_{k}^{P}-V_{k})&=\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}([Z_{k}-V_{k+1}][\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})]|\operatorname{\mathcal{F}}_{t_{k}}))+\\ &\qquad\operatorname{\mathbb{E}}[(V_{k+1}^{P}-V_{k+1})\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})]\\ &=\operatorname{\mathbb{E}}([Z_{k}-\operatorname{\mathbb{E}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})][\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})])+\\ &\qquad\operatorname{\mathbb{E}}[(V_{k+1}^{P}-V_{k+1})\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})]\\ &=\operatorname{\mathbb{E}}((Z_{k}-F_{k})[\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})])+\\ &\qquad\operatorname{\mathbb{E}}(V_{k+1}^{P}-V_{k+1})\operatorname{\mathbb{E}}(\Theta(P_{k}[V^{P}_{k+1}]-Z_{k}))+\operatorname{\mathrm{Cov}}[(V_{k+1}^{P}-V_{k+1}),\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})],\\ \end{split}$ (2.48)
@@
$\operatorname{\mathbb{E}}(|P_{k}[V^{P}_{k+1}]-F_{k})|^{2})$$\leq$ $\operatorname{\mathbb{E}}(|P_{k}[V^{P}_{k+1}]-P_{k}[V_{k+1}]|^{2})+\operatorname{\mathbb{E}}(|P_{k}[V_{k+1}]-F_{k}|^{2})$
@@
Next, we consider the asymptotic behavior of the error in the limit of $M\to\infty$ and $k\to 1$. As we can see from the fixed point analysis (Appendix A.6.2) of the backward sequence, if Assumption 2 also holds, the fixed point condition of Eqn.4.10 is satisfied with $a_{0}=\mathcal{X}^{P}_{0}$ and $c_{0}=\mathcal{Y}^{P}_{0}$. Consequently, using Eqn.4.9 and substituting $b=\Xi_{Max}^{P}$, one can see that the error saturates at
@@
Figure 2.1: The upper bound of accumulated option price error for LSM and FD-LSM at each step going backward, using the parameters mentioned in above. The $x$-axis is $M-k$ where M=60.
@@ 2.3.4 Implementation Details
where $S_{B(W)}$ represents the spot of constant weighted basket (worst of basket) with subscript $B$($W$) and we omit the time dependence on $S_{i}$ and $S_{j}$ for notational convenience.
@@
$\bar{q}$ ::=:= $-\frac{1}{T}\ln(\frac{1}{d}\sum_{i=1}^{d}e^{-q_{i}T}),$
$\bar{\sigma}^{2}$ ::=:= $\frac{1}{T}\ln\left(\frac{\frac{1}{d^{2}}\sum_{i,j}e^{(-q_{i}-q_{j}+\rho_{i,j}\sigma_{i}\sigma_{i})T}}{(\frac{1}{d}\sum_{i=1}^{d}e^{-q_{i}T})^{2}}\right).$
@@
In practice, to be used in the simulation, $f^{FD,B(W)}_{k}(x)$ is approximated as an interpolation function using a cubic spline on $V^{FD}(S,t^{+}_{k})$ along the spatial direction in the PDE grid subjected to natural boundary condition; outside the PDE grid, a linear extrapolation is used instead. In the case of worst of basket, one needs to build all individual underlying cubic splines on $\{f^{FD,i}_{k}(x)|1\leq i\leq d\}$ for a given $t_{k}$; and then for a given pathwise input $\{S_{i}(t_{k})\}$, we use the calculated $i^{*}$-th spline to do interpolation on $\min S_{i}(t_{k})$ as spatial argument.

### v10

# Finite-Difference Solution Ansatz approach in Least-Squares Monte Carlo

@@ Abstract
This article presents a simple but effective and efficient approach to improve the accuracy and stability of Least-Squares Monte Carlo. The key idea is to construct the ansatz of conditional expected continuation payoff using the finite-difference solution from one dimension, to be used in linear regression. This approach bridges between solving backward partial differential equations and Monte Carlo simulation, aiming at achieving the best of both worlds. In a general setting encompassing both local and stochastic volatility models, the ansatz is proven to act as a control variate, reducing the mean squared error, thereby leading to a reduction of the final pricing error. We illustrate the technique with realistic examples including Bermudan options, worst of issuer callable notes and expected positive exposure on European options under valuation adjustments.
Keywords: Least-squares regression, finite-difference method, Bermudan option, worst of issuer callable note, credit valuation adjustment, high-dimensional derivative pricing
@@ 1 Introduction
Solving partial differential equations (PDE) with finite-difference methods and Monte Carlo simulation are two major advanced numerical methods for exotic derivative pricing among practitioners. As the dimensionality of the problem increases, the PDE approach becomes less feasible due to implementation complexity and the numerical burden; specifically, there is no efficient and accurate implementation widely accepted in the industry with dimension greater than two. On the other hand, the Monte Carlo method is widely used in high-dimensional derivative pricing without suffering from the ‘curse of dimensionality’ as PDE with grid approach does.
@@
In this paper, we propose an enhanced least-squares regression based method by Longstaff and Schwartz, (2001), which aims at improving the projection of conditional expectations by constructing an ansatz from a finite-difference PDE solver, to be used in the linear regression. Previous studies in the literature have explored methodologies that aim to combine PDE-based and simulation-based approaches, to leverage the strengths of both. For instance, Farahany et al., (2020) introduced a method by mixing a Fourier transforms based PDE solver on discretized volatility space and regressing the value function onto these resulting conditional payoff profiles, under stochastic volatility model. However, our work is distinguished from their proposal in below aspects:
@@
The rest of this paper is organized as follows. Sec.2 is the description of the framework from a theoretical standpoint, where the underlying model settings are presented in Sec.2.1. Then we introduce background knowledge in Sec.2.2 in the context of American-style exercise derivatives pricing and CVA. After setting up the general problem in Sec.2.2.1, we outline prevailing numerical methods in practice, including the least-squares Monte Carlo approach (Sec.2.2.2 and its limitation in Sec.2.2.3), and finite-difference PDE method (Sec.2.2.4). With such foundation, and further motivated by the analysis of a limiting case in Sec.2.3.1, we proceed to introduce our novel approach in Sec.2.3, by highlighting its enhancement over the classical approach in Sec.2.3.2 and subsequently performing error analysis in Sec.2.3.3. Then implementation details will be given in Sec.2.3.4. The numerical results are delivered in Section 3, which compares the new method against existing alternatives regarding accuracy, efficiency, and stability: Sec.3.1, 3.2, and 3.3 examine Bermudan option pricing under the Black-Scholes (up to 4D) and Heston models; Sec.3.4 details CVA calculations for European options under 4D; last but not least, Sec.3.5 presents examples of worst-of issuer callable notes under a multi-variate local volatility model. Finally, concluding remarks are given in Sec.4.
@@ 2.2.1 Problem Formulation
[…]
@@
[…]
@@ 2.2.2 Least-Squares Monte Carlo
In Eqn.2.5, $\mathcal{L}(Y(\omega,t);\vec{b};\vec{\phi}^{P}_{R}(X(\omega,t_{k})))$ is the objective function to be minimized, which is the residual sum of squares (RSS). Its lower bound is given by, in the limit as $\vec{b}\cdot\vec{\phi}^{P}_{R}(X(\omega,t_{k}))\to\operatorname{\mathbb{E}}[Y(\omega,t)|X(\omega,t_{k})]$,
@@
where the exercise indicator function is defined as $I(\omega,t_{k}):=\Theta\left(\pm[Z(\omega,t_{k})-F(\omega,t_{k})]\right)$, with Heaviside step function denoted by $\Theta(x)$ and ‘++’(‘-’) for holder (issuer) exercise. The formal dynamic programming solution of $V(\omega,\tau(t_{k}))$ is also given by
@@
$\begin{split}CVA&=(1-R_{ctpy})\operatorname{\mathbb{E}}\left[\int_{0}^{T}e^{-r\cdot t}\cdot\max\left(0,\tilde{F}(\omega,t)\right)dPD(t)|{\operatorname{\mathcal{F}}}_{0}\right]\\ &=(1-R_{ctpy})\operatorname{\mathbb{E}}\left[\int_{0}^{T}e^{-r\cdot t}\cdot\max\left(0,\tilde{F}(\omega,t)\right)h_{cpty}(\omega,t)e^{-\int_{0}^{t}h_{cpty}(\omega,s)ds}dt|{\operatorname{\mathcal{F}}}_{0}\right]\\ &\approx(1-R_{ctpy})\sum_{k=1}^{M}\operatorname{\mathbb{E}}[\Phi^{CVA}_{t_{k}}|{\operatorname{\mathcal{F}}}_{0}],\end{split}$ (2.15)
@@
[…]
@@ 2.2.3 Basis function and Limitations
2.2.4 Finite-difference approach in PDE
As opposed to LSM, alternative numerical approach to solve the American-style option pricing is the finite-difference (FD) method. The elegance of this method lies in the fact that it marches from maturity to present, which naturally fits into the requirement of backward recursive optimization in stopping time problems. We restrict the discussion within 1D for maximum efficiency. The local volatility setting as describe in Sec.2.1 with $d=1$ is solvable under 1D FD PDE, under which the option price $V^{FD}(S,t)$ as function of underlying spot $S$ (omitting the subscript $i$ for the mono underlying case) and time $t$ satisfies:
@@ 2.3.2 FD Solution: Control Variate
Here, we decompose an $L^{2}$ projection using two sets of basis functions, $\vec{\phi}^{P^{\text{FD}}}(x)$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$, effectively implementing a Frisch-Waugh-Lovell (FWL) type decomposition into $P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}$ and $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$. We will focus on a detailed examination of $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$, noting that $P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}$ can be analyzed symmetrically by swapping the roles of the FD and LSM basis functions.
@@ 2.3.3 Error Analysis
[…]
@@
  • In the limit as $\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})\to 1$, the result degenerates to Lemma 1 as expected.
  • In the limit as $\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})\to 0$, the control variate has null effect as $f^{\text{FD}}_{k}(X_{k})$ is orthogonal to $F_{k}$ in $L^{2}$. In this regime, the remaining mild error reduction, going as $\sqrt{\frac{R-1}{R}}$ is purely from reducing susceptibility of overfitting by removing unnecessary high order monomial component under the constraint of Eqn.2.38; further taking the limit of $R\to\infty$, $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ converges to $P_{k}^{\text{LSM}}$ with MSE being equal to each other.
@@
in the limit as $M\to\infty$, where $\mathcal{X}^{P}_{0}:=\lim_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}\mathcal{X}^{P}_{k}$ and $\mathcal{Y}^{P}_{0}:=\lim_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}\mathcal{Y}^{P}_{k}$. For avoidance of doubt, we use “0” as the subscript to represent this limiting case where $k$ is indexed from 1 as $k\in[1,M]$.
@@
[…]
@@
$\operatorname{\mathbb{E}}(|P_{k}[V^{P}_{k+1}]-F_{k}|^{2})$ $\leq$ $\operatorname{\mathbb{E}}(|P_{k}[V^{P}_{k+1}]-P_{k}[V_{k+1}]|^{2})+\operatorname{\mathbb{E}}(|P_{k}[V_{k+1}]-F_{k}|^{2})$
@@
Next, we consider the asymptotic behavior of the error in the limit as $M\to\infty$ and $k\to 1$. As we can see from the fixed point analysis (Appendix A.6.2) of the backward sequence, if Assumption 2 also holds, the fixed point condition of Eqn.4.10 is satisfied with $a_{0}=\mathcal{X}^{P}_{0}$ and $c_{0}=\mathcal{Y}^{P}_{0}$. Consequently, using Eqn.4.9 and substituting $b=\Xi_{Max}^{P}$, one can see that the error saturates at
@@
Figure 2.1: The upper bound of accumulated option price error for LSM and FD-LSM at each step going backward, using the parameters mentioned in above. The $x$-axis is MkM-k where M=60.
@@ 2.3.4 Implementation Details
where $S_{B(W)}$ represents the spot of constant weighted basket (worst of basket) with subscript BB($W$) and we omit the time dependence on $S_{i}$ and $S_{j}$ for notational convenience.
@@
$\bar{q}$ $:=$ $-\frac{1}{T}\ln(\frac{1}{d}\sum_{i=1}^{d}e^{-q_{i}T}),$
$\bar{\sigma}^{2}$ $:=$ $\frac{1}{T}\ln\left(\frac{\frac{1}{d^{2}}\sum_{i,j}e^{(-q_{i}-q_{j}+\rho_{i,j}\sigma_{i}\sigma_{i})T}}{(\frac{1}{d}\sum_{i=1}^{d}e^{-q_{i}T})^{2}}\right).$
@@
[…]